STAR Protected
February 23
- RAW to MuDst production
- pp500_production_2017 (physics)
409176 files have been converted out of 654286
Around 64% Completion
- pp500_production_2017 (physics)
CVS->Git
On June 1, 2021, the main repository for STAR offline software has migrated from CVS to GitHub (https://github.com/star-bnl).
February 16
- RAW to MuDst production
- pp500_production_2017 (physics)
394786 files have been converted out of 654286
Around 61% Completion
- pp500_production_2017 (physics)
TPC calibration reviews
virtual , at 14:30 (GMT), duration : 00:00
https://indico.bnl.gov/event/10740/
This page will be used for draft talks
Time | Talk | Presenter |
---|---|---|
09:00 | TPC Calibration Organization ( 00:20 ) 1 file | Gene Van Buren (BNL) |
09:30 | surveys ( 00:20 ) 1 file | F.Videbaek |
09:50 | Resources for HLT and Express Production ( 00:20 ) 1 file | Hongwei Ke |
11:00 | TPC Distortions ( 00:40 ) 1 file | Gene Van Buren (BNL) |
11:40 | FXT event-by-event correction ( 00:20 ) 1 file | Irakli |
12:00 | TPC calibration ( 01:10 ) 1 file | Yuri Fisyak |
February 9
- RAW to MuDst production
- pp500_production_2017 (physics)
394672 files have been converted out of 654286
Around 61% Completion
- pp500_production_2017 (physics)
February 2
- RAW to MuDst production
- pp500_production_2017 (physics)
394469 files have been converted out of 654286
Around 60% Completion
- pp500_production_2017 (physics)
New Gating Grid Driver issues
8/4/2021
The new GG have been setup and calibrated, and is getting ready for tests.
TPC weekly
, at 20:00 (GMT), duration : 00:00
agenda
- hardware status
- software updates
- calibration review
TPC weeklu
, at 20:00 (GMT), duration : 00:00
TPC weekly
, at 20:00 (GMT), duration : 00:00
hardware
- the west laser had a LV power supply with a bad fuse. Alexie got an access fand fixed in < 1 hour
- Alexei monitor gas and will adjust methan concentration slightly.
- new hot channel in sector1. Have asked Tonko to mask out
FF request.
Gene summarized the FF request as it had emerged from the calibration review. It is important to run FF setting to seperate alignment, E*B distortins
but having data from both polarities. Yuri & Gene confirm that the 20M RF is a good sample from RFF and it was concluded we should request a change as soon as practical and on Wednesday access day at latest. Alexei brought up request to at at the time of switch that a high stat laser run ~20k is taken jest before magnet change and another one is taken with FF after the change, and magnet stabilized.
Yuri showed the dedx for his run 20 calibration for collider data and fixed target.There is some room for improvement in particular for the FXT target data.
Yuri showed that the readout of the magnets current fluctuate on a 15 sec time sacle between two distinct values. Current code only has one magnet value per run so it gives rise to distinct bands of say K0_as there is a scale factor between current read and Field suggest to use average value for future.
Gene discussed a Minor problem in sector alignment distortion correction as shown in attached slide on web page.
Looked at run 18 distortion maps and notice a missing featue that he expected namely spikes missing between inner and outer sector. This is due to an un intentional change of code in 2014. Demonstrated with differences that these effect are likely small. One measure of #primary tracks is sub-promille.
Time | Talk | Presenter |
---|---|---|
15:00 | Run XX dE/dx calibration ( 00:05 ) 0 files | Yuri Fisyak |
15:05 | Run XXI Ks0 mass and STAR main magnet current ( 00:05 ) 0 files | Yuri Fisyak |
15:10 | Minor problem in sector alignment distortion correction ( 00:20 ) 0 files | G. Van Buren (BNL) |
TPC weeklu
, at 20:00 (GMT), duration : 00:00
TPC weekly
, at 20:00 (GMT), duration : 00:00
- hardware status
- run progress
- software
- TPC Review Planning
Charge to committee
We just formed the STAR TPC calibration review committee. It is timely and critical to look at the offline calibration tasks and issues. The committee is charged to evaluate the readiness of the TPC calibration for the BES-II data production and provide answers to the following questions by the end of March 2021:
1. What is the status of the tracking and production calibration?
2. Where is the bottleneck of the TPC calibration procedure?
3. Are there enough collaboration manpower and resources for the tasks?
4. How are the online/operation/fast-offline/PWGs efforts coordinated to accomplish the goals?
5. Is there a clear document of procedures in calibration and what are the unexpected and expected issues in BES-II datasets? What are the steps in calibration, and how are the current responsibilities assigned?
6. What is the timeline and path toward completion of whole BES-II calibration for production?
7. Any lesson learned toward future runs?
The members are:
Zhangbu Xu (Chair)
Bill Christie
Carl Gagliardi
Frank Geurts
Marian Ivanov
Ivan Kisel
Tonko Ljubicic
Maria Stefaniak
Jim Thomas
Qinghua Xu
Xin Dong (ex officio)
Rongrong Ma (ex officio)
Talks and subject area.
Geometry & alignment (Yuri)
• Drift velocities
• dE/dx
• Specifics for Run 19
• Cosmic data
• Beam data
• Laser data
• What is the procedure for alignment; How is it documented
• Bottleneck in timely calibrations
• Validation procedures and QA
• Lessons and procedures learned from iTPC prototype commissioning
• Present/summarize clearly the open/unknown issues and potential concerns (all talks)
Surveys done (Flemming)
• Alexei, Jim and Yuri (operation survey)
• Other detectors as calibration tools (GMT, TOF, BEMC)
• How they are included in the calibration procedure
Correction Distortions (Gene)
• Dynamic distortions: SpaceCharge & GridLeak
• Static distortions
Resources for Online and express stream (Hongwei)
• Recap of resources
• Activities
• Relation to FastOffline
Calibration (Irakli)
• T0
• Long bunch timing, FXT and collider BES-II
• How is express and fast offline used in calibration
• Resource/tools/datasets use
QA,QC (BES-II) (Dan)
• Recap of resources
• How are working groups involved in this
• Feedback, requirements to calibrations
TPC weekly
, at 20:00 (GMT), duration : 00:00
-- hardware status
-- startup plans
electronics update (Tonko)
a brief update regarding both TPX (aka outer sectors)
and iTPC (inner sectors) electronics and gating grid.
1) We have the new (Tim's) gating grid connected.
I calibrated it and it looks OK to me. But we should be
a bit weary because we had problems before.
One outcome is that it is relatively noisier at the
opening time than the old GG so one should pay attention.
2) The pulser runs have been taken and the gains
have been established as well as the bad-channel
list. Pretty much the same as last year because
we had no time (COVID and all) to mess with swapping
FEEs to check the problem and see where the
noise comes from: FEE or padplane.
The gains have been loaded into the database.
3) We _did_ swap the broken FEEs and RDOs.
As luck would have it, a single FEE in the ITPC worked
all Fall/Winter but failed as we closed the poletip.
It is masked.
Let's take it from here and see how we stand during
Charge for calibration review
Dear All,
We just formed the STAR TPC calibration review committee. It is timely and critical to look at the offline calibration tasks and issues. The committee is charged to evaluate the readiness of the TPC calibration for the BES-II data production and provide answers to the following questions by the end of March 2021:
- What is the status of the tracking and production calibration?
- Where is the bottleneck of the TPC calibration procedure?
- Are there enough collaboration manpower and resources for the tasks?
- How are the online/operation/fast-offline/PWGs efforts coordinated to accomplish the goals?
- Is there a clear document of procedures in calibration and what are the unexpected and expected issues in BES-II datasets? What are the steps in calibration, and how are the current responsibilities assigned?
- What is the timeline and path toward completion of whole BES-II calibration for production?
- Any lesson learned toward future runs?
The members are:
Zhangbu Xu (Chair)
Bill Christie
Carl Gagliardi
Frank Geurts
Marian Ivanov
Ivan Kisel
Tonko Ljubicic
Maria Stefaniak
Jim Thomas
Qinghua Xu
Xin Dong (ex officio)
Rongrong Ma (ex officio)
Time | Talk | Presenter |
---|---|---|
15:00 | Residuals re-tested ( 00:20 ) 0 files | Irakli |
15:20 | Comparison express production TFG19e and calibration production P20ic_calib for 19GeV AuAu2019 data ( 00:20 ) 0 files |
January 26
- RAW to MuDst production
- pp500_production_2017 (physics)
394125 files have been converted out of 654286
Around 60% Completion
- pp500_production_2017 (physics)